A social psychologist wants to test if extroversion is significantly related to a person’s general health. He gets extroversion scores and a measure of general health from the participants’ medical providers. The data are listed below. Higher health extroversion scores mean greater extroversion, while high health scores = good health.  

Complete the appropriate statistical test and hypothesis testing using the 4 step approach. State your conclusion in APA format. Alpha level = .05.

Participant

extrov X

health Y

X squared Y squared (X) (Y)
A

101

33 3333
B 93 40 3720
C 85 24 2040
D 111 46 5106
E 104 30 3120
F 89 28 2492
Total
Answer #1

From the given information, xy 24 Participant extrov (x) health (y) x? y A 101 33 10201 1089 3333 B 93 40 8649 1600 3720 C 85Step-1: Null hypothesis, H.: there is no relationship between IQ score related and to health Alternative hypothesis, H, : thedf =n-2 = 6-2 = 4 Step-4: Calculate the P-value. (Use MS Excel function) P-value =(=TDIST(1.96,4,2)) = 0.121555 Decision: Fai